OpenStudy (anonymous):

What are the exact solutions of x^2 − 3x − 1 = 0?

OpenStudy (anonymous):

I got two choices I found, but not sure which is the right one.

OpenStudy (anonymous):

Use the quadratic formula.

OpenStudy (anonymous):

I think both solutions are the right one because quadratic equations have 2 solutions

OpenStudy (anonymous):

\[x = \frac{ -3 \pm \sqrt{13} }{ 2 }\] \[x = \frac{ 3 \pm \sqrt{13} }{ 2 }\]

OpenStudy (anonymous):

I have to pick one.

OpenStudy (anonymous):

Oh lol 4 solutions

OpenStudy (anonymous):

The formula starts with -b and b = -3 so -(-3) = 3, the second set is the right one

OpenStudy (anonymous):

Maybe I need to listen to my teacher more because she said that same exact thing and forgot about it. lol.